# AdapterSwap: Continuous Training of LLMs with Data Removal and Access-Control Guarantees

### v1

@@ 5.5 Data Removal
We also measure our ability to efficiently purge documents from our dataset after training. When a piece of data is removed from the corpus, only the adapter fine-tuned on that data requires retraining.
@@
Table 4: Perplexity completing documents using an adapter trained with (Before Purge) and without (Purged) the purged data.
We show this by attempting to complete documents for each adapter before and after removing them and retraining the adapter.
@@ 6.1 Knowledge Editing
Recent work in knowledge editing of LLMs has considered approaches which either add additional parameters to a model, or directly edit existing parameters to update information (Yao et al., 2023). Directly updating the existing parameters is attractive as it does not require any additional parameters, and updates can be applied whenever new knowledge is available (De Cao et al., 2021; Meng et al., 2022). Knowledge vectors, in combination with hidden representations of specific entities, have also been proposed as a tool to update or remove knowledge (Hernandez et al., 2023). In all cases, these approaches lack the ability to guarantee that any specific training example can be removed entirely from the model.
@@ 6.2 Retrieval Augmented Generation
Retrieval-Augmented Generation (RAG) has become a popular approach to incorporating new data into a pre-trained LLM without having to rely on retraining or fine-tuning (Lewis et al., 2021; Gao et al., 2024). While it offers some advantages, there remain challenges that can make deploying an effective RAG-based solution difficult. For example, RAG is limited by how much retrieved context can be employed based on the underlying LLM’s context window size. This contrasts with AdapterSwap where each adapter in our experiments represented more than 50 million tokens on average. Recent work has shown that all evidence in the context window is not treated equally, with models favoring evidence at the start and end of each window (Liu et al., 2023). Further, RAG is fully dependent on the ability of a retriever model to locate all relevant documents without introducing too much noise into the context (Barnett et al., 2024; Gao et al., 2024; Reyes, 2024). In addition, as transformers are quadratic in the number of tokens considered, then requiring forced decoding over retrieved content can add significant latency at inference time. While our solution avoids these issues, we note that AdapterSwap does not preclude the use of RAG, and a hybrid approach could be useful in some circumstances.
@@ 6.3 Federated Learning
Federated learning (McMahan et al., 2017) also deals with learning from siloed data, typically aggregating gradients on local data before averaging into a global model. However, federated learning is intended to produce a single centralized model without mixing data silos and thus does not provide any mechanism for access control or deletion. Some work has combined federated learning and PEFT methods (Kim et al., 2023; Babakniya et al., 2023; Zhang et al., 2023), but these do not address adapter mixing, deletion, or data silos as distinct knowledge sources. Furthermore, the privacy benefits of federated learning are unclear when applied to LLMs with large capacity for memorization (Gupta et al., 2022; Shi et al., 2023; Tirumala et al., 2022; Carlini et al., 2023).
7 Conclusion and Future Work
@@ Appendix A Training Details
For all experiments we initialized adapters using a random seed of 42. We confirm Chronopoulou et al. (2023)’s finding that using the same initialization is critical if mixing adapters.

### v2

@@ 5.5 Data Removal
We also measure our ability to efficiently purge documents from our dataset after training. When a piece of data is removed from the corpus, only the adapter fine-tuned on that data requires retraining. We show this by attempting to complete documents for each adapter before and after removing them and retraining the adapter.
Table 4: Perplexity completing documents using an adapter with (Before Purge) and without (Purged) the purged data.
@@ 6.1 Knowledge Editing
Recent work in knowledge editing of LLMs has considered approaches which either add additional parameters to a model, or directly edit existing parameters to update information [35]. Directly updating the existing parameters is attractive as it does not require any additional parameters, and updates can be applied whenever new knowledge is available [36, 37]. Knowledge vectors, in combination with hidden representations of specific entities, have also been proposed as a tool to update or remove knowledge [38]. In all cases, these approaches lack the ability to guarantee that any specific training example can be removed entirely from the model.
@@ 6.2 Retrieval Augmented Generation
Retrieval-Augmented Generation (RAG) has become a popular approach to incorporating new data into a pre-trained LLM without having to rely on retraining or fine-tuning [39, 40]. While it offers some advantages, there remain challenges that can make deploying an effective RAG-based solution difficult. For example, RAG is limited by how much retrieved context can be employed based on the underlying LLM’s context window size. This contrasts with AdapterSwap where each adapter in our experiments represented more than 50 million tokens on average. Recent work has shown that all evidence in the context window is not treated equally, with models favoring evidence at the start and end of each window [41]. Further, RAG is fully dependent on the ability of a retriever model to locate all relevant documents without introducing too much noise into the context [42, 40, 43]. In addition, as transformers are quadratic in the number of tokens considered, then requiring forced decoding over retrieved content can add significant latency at inference time. While our solution avoids these issues, we note that AdapterSwap does not preclude the use of RAG, and a hybrid approach could be useful in some circumstances.
@@ 6.3 Federated Learning
Federated learning [44] also deals with learning from siloed data, typically aggregating gradients on local data before averaging into a global model. However, federated learning is intended to produce a single centralized model without mixing data silos and thus does not provide any mechanism for access control or deletion. Some work has combined federated learning and PEFT methods [45, 46, 47], but these do not address adapter mixing, deletion, or data silos as distinct knowledge sources. Furthermore, the privacy benefits of federated learning are unclear when applied to LLMs with large capacity for memorization [48, 49, 50, 51].
7 Conclusion
@@ Appendix A Training Details
  • For all experiments we initialized adapters using a random seed of 42. We confirm [21]’s finding that using the same initialization is critical if mixing adapters.